A 50.0g thermometer is used to measure the temperature of 170 mL of water. The specific heat of the thermometer, which is mostly glass, is 750J/(kgK), and it reads 18.0 ∘C while lying on the table. After being completely immersed in the water, the thermometer's reading stabilizes at 75.6 ∘C. Part A What was the actual water temperature before it was measured?
Answer #1

Energy released = energy absorbed

Q = m*Cp*dT

Mw = 170*10^-3 L = 0.17 kg

Mt = 50.0 gm = 0.05 kg

Cw = 4186 J/kg-C

Cc = 750 J/kg-C

So,

Qw = Qc

0.17*4186*(T - 75.6) = 0.05*750*(75.6 - 18.0)

Solving the equation:

T = 78.63 C